OpenStudy (anonymous):

Can someone please check this for me? There not hard at all..but to me, of course they are -_-"

jhonyy9 (jhonyy9):

y=mx +b formula is hence yes ?

jhonyy9 (jhonyy9):

how do you think if you grafh this with examples i think you will get the right answer sure easy ,yes ?

jhonyy9 (jhonyy9):

how is your equation right ?can you rewrite it here now please ?

jhonyy9 (jhonyy9):

first y=2^x ,right ? so than in this case when x=0 result y=1 and when x=1 result y=2 grafh this please and after calcule how many will result for y value when y= 2^(x-3) +4 so in the same like before you add value to x and see what will get than for y grafh these values and see what will be the difference from this grafh and the firstly grafh than will get the right anser to the first question easy ok ?

jhonyy9 (jhonyy9):

sorry yes there is k and not x

jhonyy9 (jhonyy9):

no problems good luck bye